### v3

@@ 3 Lyapunov and Riccati inequalities
with ${\bf Y}=\hat{\bf S}^{-1}({\bf B}^{\top}{\bf W}_{2}-{\bf C}{\bf W}_{1})$, i.e., the columns of $\left[\begin{array}[]{ccc}-{\bf W}_{2}^{\top}&{\bf W}_{1}^{\top}&{\bf Y}^{\top}\end{array}\right]^{\top}$ span the $n$-dimensional deflating subspace of $\lambda\boldsymbol{\mathcal{N}}-\boldsymbol{\mathcal{M}}$ corresponding to the finite eigenvalues contained in ${\bf Z}$.
A pencil $\lambda\,\boldsymbol{\mathcal{N}}-\boldsymbol{\mathcal{M}}$ is called even if $\boldsymbol{\mathcal{N}}=-\boldsymbol{\mathcal{N}}^{\top}$ and $\boldsymbol{\mathcal{M}}=\boldsymbol{\mathcal{M}}^{\top}$. Since even pencils have the Hamiltonian spectral symmetry in the finite eigenvalues, see ByeMX07 , this means that there are equally many eigenvalues in the open left and in the open right half plane.
Numerically, to compute ${\bf Q}(={\bf W}_{2}{\bf W}_{1}^{-1})$ it is preferable to work with the even pencil (59) rather than with the Hamiltonian matrix (43), since explicit inversion of $\hat{\bf S}$ is avoided. Numerically stable structure preserving methods for computing an orthonormal matrix $\left[\begin{array}[]{ccc}-{\bf W}_{2}^{\top}&{\bf W}_{1}^{\top}&{\bf Y}^{\top}\end{array}\right]^{\top}$ are available, see BenBMX99 ; BenLMV15 ; BruM08 .
Remark 20.
@@
Secondly, for (19) or (20), we still require that ${\bf A}_{11}-{\bf B}_{1}\hat{\bf S}^{-1}{\bf C}_{1}$ is asymptotically stable and that the Hamiltonian matrix ${\bf H}_{11}$ in (48) has a Lagrangian invariant subspace. Since this only requires that ${\bf Q}_{1}\geq 0$, a solution can be determined in a simpler way. We may simply set $\widetilde{\bf Q}_{11}=\widetilde{\bf Q}_{11}^{0}$ and $\widetilde{\Psi}_{11}=0$. Then with the block structure the solution of $\widetilde{\Psi}_{12}=0$ has a form $\widetilde{\bf Q}_{12}=\left[\begin{array}[]{c}{\bf Q}_{12}\\ 0\end{array}\right]$. To solve $\widetilde{\Psi}_{22}=0$ for $\widetilde{\bf Q}_{22}$, one can show ${\bf Q}_{1}={\bf U}\left[\begin{array}[]{cc}\widetilde{\bf Q}_{11}&\widetilde{\bf Q}_{12}\\ \widetilde{\bf Q}_{12}^{\top}&\widetilde{\bf Q}_{22}\end{array}\right]{\bf U}^{\top}\geq 0$ and solves the Riccati equation $\Psi({\bf Q}_{1})=0$. Also, in some circumstances the condition that $\widetilde{\bf A}_{22}={\bf A}_{33}$ is asymptotically stable can be relaxed. In these relaxed cases however, it is necessary to transform the system to a descriptor formulation, see CheGH22 .
@@ 4 Numerical procedures for the construction of port-Hamiltonian realizations
The procedures presented in the following three subsections follow from the procedures in WeiWS94 . The only extra result is an explicit formulation of the part of the matrix ${\bf T}$ for dealing with the singularity of ${\bf D}+{\bf D}^{\top}$. We provide proofs for completeness.
@@
Suppose that the matrix $\hat{\bf S}={\bf D}+{\bf D}^{\top}\geq 0$ is singular. Consider an orthogonal matrix ${\bf V}_{0}=[{\bf V}_{0,1},{\bf V}_{0,2}]$, where ${\bf V}_{0,1}$ is chosen so that its columns form an orthonormal basis of the kernel of $\hat{\bf S}$. To construct such a ${\bf V}_{0}$ we can use a singular value or rank-revealing $QR$ decomposition, GolV96 . Then we have
@@
where ${\bf S}_{2}={\bf S}_{2}^{\top}>0$. Set
@@
Remark 21.
If $\hat{\bf S}$ is positive definite but nearly singular, then the coefficients of the Riccati inequality in (17) may incur large relative errors when formed. In this case, it may be appropriate to regularize the problem by perturbing $\hat{\bf S}$ to a nearby positive semidefinite (but now singular) problem as in (62) where the matrix ${\bf S}_{2}$ is now well conditioned with respect to inversion. In our procedure we do this by setting small positive eigenvalues of $\hat{\bf S}$ to zero.
@@
has an invertible solution ${\bf T}$ satisfying the constraint $({\bf T}{\bf B}_{1})^{\top}-{\bf C}_{1}{\bf T}^{-1}=0$. We characterize conditions when this constraint is satisfied in the following subsections. But notice that we are, in effect, restricting the input and output space to the invertible part of ${\bf D}+{\bf D}^{\top}$. Once these restricted transformation matrices have been constructed, full transformations satisfying the given constraint can be obtained by extending to the full space.
@@ 4.2 Construction in the case ${\bf D}=-{\bf D}^{\top}$
Note that without the constraint (66a), if ${\bf A}$ is stable, then by Lemma 9 the second condition (66b) can always be satisfied. Adding the constraint (66a), however, makes the question nontrivial.
@@
Lemma 23.
WeiWS94 Consider ${\bf B},{\bf C}^{\top}\in{\mathbb{R}}^{n\times m}$, and assume that $\mathop{\rm rank}\nolimits{\bf B}=r$.
@@
Proof.
Condition (66a) is equivalent to ${\bf C}={\bf B}^{\top}{\bf T}^{\top}{\bf T}$ and the following conditions will be necessary for the existence of an invertible ${\bf T}$ with this property:
@@
with ${\bf B}_{1},{\bf C}_{1}^{\top}\in{\mathbb{R}}^{n\times r}$ having full column rank. The conditions $\mathop{\rm rank}\nolimits{\bf C}{\bf B}=r$ and ${\bf C}{\bf B}\geq 0$ together are equivalent to $\widetilde{\bf C}_{1}\widetilde{\bf B}_{1}>0$. Thus, there must exist an invertible matrix ${\bf Y}$ (e.g., a Cholesky factor or a positive-definite square root) such that ${\bf C}_{1}{\bf B}_{1}={\bf Y}{\bf Y}^{\top}$.
@@
In order to explicitly construct a transformation matrix ${\bf T}_{0}$ as in (67), it will be useful to construct bi-orthogonal bases for the two subspaces $\mathsf{Ker}\,{\bf B}^{\top}$ and $\mathsf{Ker}\,{\bf C}$. Toward this end, let ${\bf N}_{{\bf C}}$ contain columns that form a basis of $\mathsf{Ker}\,{\bf C}$, so that $\mathsf{Ran}\,{\bf N}_{{\bf C}}=\mathsf{Ker}\,{\bf C}$. Such a matrix is easily constructed in a numerically stable way via the singular value decomposition or a rank-revealing $QR$ decomposition of ${\bf C}$, see GolV96 . Since ${\bf B}$ and ${\bf C}^{\top}$ are assumed to satisfy $\mathsf{Ker}\,{\bf B}=\mathsf{Ker}\,{\bf C}^{\top}$, we have singular value decompositions
@@
Proof.
The condition follows from Corollary 22 and the representation (67) by setting ${\bf U}={\bf I}$ and ${\bf T}_{0}$ as in (75).
@@ 4.3 Construction in the case of general ${\bf D}$
For the case that ${\bf D}$ is general we will present a recursive procedure which is analogous to the index reduction procedure for differential-algebraic equations in KunM24 . The first step is to perform the transformations (62), (63), and to obtain the following characterization when a transformation to port-Hamiltonian form (6) exists.
Lemma 25.
Consider system (5) transformed as in (62) and (63). Then the system is equivalent to a port-Hamiltonian system of the form  (6) if and only if
@@
    $\mathsf{Ker}\,{\bf C}_{1}^{\top}=\mathsf{Ker}\,{\bf B}_{1}$, $\mathop{\rm rank}\nolimits{\bf C}_{1}{\bf B}_{1}=\mathop{\rm rank}\nolimits{\bf B}_{1}$, and ${\bf C}_{1}{\bf B}_{1}\geq 0$, or equivalently, there exists an invertible (orthogonal) matrix ${\bf W}$ such that
@@
    $\widetilde{\bf Y}:=\left[\begin{array}[]{cc|c}{\bf Z}&0&0\\ 0&{\bf I}&0\\ \hline\cr 0&0&{\bf I}\end{array}\right]\left[\begin{array}[]{c|c}-{\bf T}_{0}{\bf A}{\bf T}_{0}^{-1}&-{\bf T}_{0}{\bf B}_{2}\\ \hline\cr{\bf C}_{2}{\bf T}_{0}^{-1}&{\bf V}_{0,2}^{\top}{\bf D}{\bf V}_{0,2}\end{array}\right]\left[\begin{array}[]{cc|c}{\bf Z}&0&0\\ 0&{\bf I}&0\\ \hline\cr 0&0&{\bf I}\end{array}\right]^{-1}$ (79)
@@
    and the columns of full rank matrices $\widetilde{\bf N}_{{\bf B}_{1}}$ and $\widetilde{\bf N}_{{\bf C}_{1}}$ form the kernels of ${\bf B}_{1}^{\top}$, ${\bf C}_{1}$ respectively, and satisfy $\widetilde{\bf N}_{{\bf B}_{1}}^{\top}\widetilde{\bf N}_{{\bf C}_{1}}={\bf I}$.
Proof.
Condition (66a) in this case has the form $({\bf T}{\bf B}_{1})^{\top}={\bf C}_{1}{\bf T}^{-1}$. If this condition holds, then applying the result of Lemma 23 to ${\bf B}_{1}$ and ${\bf C}_{1}$ one has ${\bf T}$ as the one in Lemma 23 b) with ${\bf T}_{0}$ of the form as (67). The result is proved by applying this formula to (61).
We can repartition the middle factor of $\widetilde{\bf Y}$ in (79) as
$\left[\begin{array}[]{cc}-{\bf T}_{0}{\bf A}{\bf T}_{0}^{-1}&-{\bf T}_{0}{\bf B}_{2}\\ {\bf C}_{2}{\bf T}_{0}^{-1}&{\bf V}_{0,2}^{\top}{\bf D}{\bf V}_{0,2}\end{array}\right]$ $=$ $\left[\begin{array}[]{c|cc}-\widetilde{\bf N}_{{\bf B}_{1}}^{\top}{\bf A}\widetilde{\bf N}_{{\bf C}_{1}}&-\widetilde{\bf N}_{{\bf B}_{1}}^{\top}{\bf A}\widehat{\bf B}_{1}{\bf Y}^{-\top}&-\widetilde{\bf N}_{{\bf B}_{1}}^{\top}{\bf B}_{2}\\ \hline\cr-{\bf Y}^{-1}\widehat{\bf C}_{1}{\bf A}\widetilde{\bf N}_{{\bf C}_{1}}&-{\bf Y}^{-1}\widehat{\bf C}_{1}{\bf A}\widehat{\bf B}_{1}{\bf Y}^{-\top}&-{\bf Y}^{-1}\widehat{\bf C}_{1}{\bf B}_{2}\\ {\bf C}_{2}\widetilde{\bf N}_{{\bf C}_{1}}&{\bf C}_{2}\widehat{\bf B}_{1}{\bf Y}^{-\top}&{\bf V}_{0,2}^{\top}{\bf D}{\bf V}_{0,2}\end{array}\right]$ (86)
=:absent:=:= : $\left[\begin{array}[]{rr}-\widetilde{\bf A}_{1}&-\widetilde{\bf B}_{1}\\ \widetilde{\bf C}_{1}&\widetilde{\bf D}_{1}\end{array}\right].$ (89)
Thus, we have that
$\widetilde{\bf Y}=\left[\begin{array}[]{cc}{\bf Z}&0\\ 0&{\bf I}\end{array}\right]\left[\begin{array}[]{rr}-\widetilde{\bf A}_{1}&-\widetilde{\bf B}_{1}\\ \widetilde{\bf C}_{1}&\widetilde{\bf D}_{1}\end{array}\right]\left[\begin{array}[]{cc}{\bf Z}&0\\ 0&{\bf I}\end{array}\right]^{-1},$
@@
This implies that the procedure of checking the existence of a transformation from (5) to a pH system can be performed in a recursive way. One first performs the transformation (62) and checks whether a condition as in Part 1 of Lemma 25 does not hold, in which case a transformation does not exist. Otherwise, one checks whether $\widetilde{\bf D}_{1}+\widetilde{\bf D}_{1}^{\top}$ in (89) is invertible. If it is and if the associated matrix inequality does not have a positive definite solution, then a transformation does not exist. Otherwise, a transformation exists and a corresponding transformation matrix ${\bf T}$ can be constructed by computing ${\bf Z}$ satisfying $\widetilde{\bf Y}({\bf Z})+\widetilde{\bf Y}({\bf Z})^{\top}\geq 0$ and the matrix ${\bf T}_{0}$ is formed accordingly. In the remaining case, i.e., a condition as in Part 1 of Lemma 25 holds and $\widetilde{\bf D}_{1}+\widetilde{\bf D}_{1}^{\top}$ is singular, the process is repeated.
@@
$\widetilde{\bf T}_{0}=\left[\begin{array}[]{cc}{\bf T}_{0}&0\\ 0&{\bf I}\end{array}\right],\quad\widetilde{\bf V}_{0}=\left[\begin{array}[]{cc}{\bf I}&0\\ 0&{\bf V}_{0}\end{array}\right],$
@@
${\bf G}_{1}=\left[\begin{array}[]{c|cc|c}-\widetilde{\bf N}_{{\bf B}_{1}}^{\top}{\bf A}\widetilde{\bf N}_{{\bf C}_{1}}&-\widetilde{\bf N}_{{\bf B}_{1}}^{\top}{\bf A}\widehat{\bf B}_{1}{\bf Y}^{-\top}&-\widetilde{\bf N}_{{\bf B}_{1}}^{\top}{\bf B}_{2}&0\\ \hline\cr-{\bf Y}^{-1}\widehat{\bf C}_{1}{\bf A}\widetilde{\bf N}_{{\bf C}_{1}}&-{\bf Y}^{-1}\widehat{\bf C}_{1}{\bf A}\widehat{\bf B}_{1}{\bf Y}^{-\top}&-{\bf Y}^{-1}\widehat{\bf C}_{1}{\bf B}_{2}&-\widetilde{\Gamma}^{\top}\\ {\bf C}_{2}\widetilde{\bf N}_{{\bf C}_{1}}&{\bf C}_{2}\widehat{\bf B}_{1}{\bf Y}^{-\top}&{\bf V}_{0,2}^{\top}{\bf D}{\bf V}_{0,2}&{\bf V}_{0,2}^{\top}{\bf D}{\bf V}_{0,1}\\ \hline\cr 0&\widetilde{\Gamma}&{\bf V}_{0,1}^{\top}{\bf D}{\bf V}_{0,2}&{\bf V}_{0,1}^{\top}{\bf D}{\bf V}_{0,1}\end{array}\right],$
@@
${\bf C}_{1}{\bf T}_{0}^{-1}$ $=$ $\left[\begin{array}[]{cc}0&{\bf W}^{-\top}\left[\begin{array}[]{c}\widehat{\bf C}_{1}\\ 0\end{array}\right]\widehat{\bf B}_{1}{\bf Y}^{-\top}\end{array}\right]=\left[\begin{array}[]{cc}0&{\bf W}^{-\top}\left[\begin{array}[]{c}{\bf Y}\\ 0\end{array}\right]\end{array}\right]=\left[\begin{array}[]{cc}0&\widetilde{\Gamma}\end{array}\right],$
${\bf T}_{0}{\bf B}_{1}$ $=$ $\left[\begin{array}[]{c}0\\ {\bf Y}^{-1}\widehat{\bf C}_{1}\left[\begin{array}[]{cc}\widehat{\bf B}_{1}&0\end{array}\right]{\bf W}^{-1}\end{array}\right]=\left[\begin{array}[]{c}0\\ \left[\begin{array}[]{cc}{\bf Y}^{\top}&0\end{array}\right]{\bf W}^{-1}\end{array}\right]=\left[\begin{array}[]{c}0\\ \widetilde{\Gamma}^{\top}\end{array}\right].$

### v4

@@ 3 Lyapunov and Riccati inequalities
[…]
Numerically, to compute ${\bf Q}(={\bf W}_{2}{\bf W}_{1}^{-1})$ it is preferable to work with the even pencil (59) rather than with the Hamiltonian matrix (43), since explicit inversion of $\hat{\bf S}$ is avoided. Numerically stable structure preserving methods for computing an orthonormal matrix $\left[\begin{array}[]{ccc}-{\bf W}_{2}^{\top}&{\bf W}_{1}^{\top}&{\bf Y}^{\top}\end{array}\right]^{\top}$ are available, see [9, 10, 13].
Remark 20
@@
Secondly, for (19) or (20), we still require that ${\bf A}_{11}-{\bf B}_{1}\hat{\bf S}^{-1}{\bf C}_{1}$ is asymptotically stable and that the Hamiltonian matrix ${\bf H}_{11}$ in (48) has a Lagrangian invariant subspace. Since this only requires that ${\bf Q}_{1}\geq 0$, a solution can be determined in a simpler way. We may simply set $\widetilde{\bf Q}_{11}=\widetilde{\bf Q}_{11}^{0}$ and $\widetilde{\Psi}_{11}=0$. Then with the block structure the solution of $\widetilde{\Psi}_{12}=0$ has a form $\widetilde{\bf Q}_{12}=\left[\begin{array}[]{c}{\bf Q}_{12}\\ 0\end{array}\right]$. To solve $\widetilde{\Psi}_{22}=0$ for $\widetilde{\bf Q}_{22}$, one can show ${\bf Q}_{1}={\bf U}\left[\begin{array}[]{cc}\widetilde{\bf Q}_{11}&\widetilde{\bf Q}_{12}\\ \widetilde{\bf Q}_{12}^{\top}&\widetilde{\bf Q}_{22}\end{array}\right]{\bf U}^{\top}\geq 0$ and solves the Riccati equation $\Psi({\bf Q}_{1})=0$. Also, in some circumstances the condition that $\widetilde{\bf A}_{22}={\bf A}_{33}$ is asymptotically stable can be relaxed. In these relaxed cases however, it is necessary to transform the system to a descriptor formulation, see [18].
@@ 4 Numerical procedures for the construction of port-Hamiltonian realizations
The procedures presented in the following three subsections follow from the procedures in [63]. The only extra result is an explicit formulation of the part of the matrix $\bf T$} for dealing with the singularity of ${\bf D}+{\bf D}^{\top}$. We provide proofs for completeness.
@@
Suppose that the matrix $\hat{\bf S}={\bf D}+{\bf D}^{\top}\geq 0$ is singular. Consider an orthogonal matrix ${\bf V}_{0}=[{\bf V}_{0,1},{\bf V}_{0,2}]$, where ${\bf V}_{0,1}$ is chosen so that its columns form an orthonormal basis of the kernel of $\hat{\bf S}$. To construct such a ${\bf V}_{0}$ we can use a singular value or rank-revealing $QR$ decomposition, [25]. Then we have
@@
where $0<{\bf S}_{2}={\bf D}_{2}+{\bf D}_{2}^{\top}$, ${\bf D}_{2}={\bf V}_{0,2}^{\top}{\bf D}{\bf V}_{0,2}\in\mathbb{R}^{s\times s}$ and $s=\mathop{\rm rank}\nolimits({\bf D}+{\bf D}^{\top})$. Set
@@
has an invertible solution $\bf T$} satisfying the constraint $({\bf T}{\bf B}_{1})^{\top}-{\bf C}_{1}{\bf T}^{-1}=0$. We characterize conditions when this constraint is satisfied in the following subsections. Notice that we are, in effect, restricting the input and output space to the invertible part of ${\bf D}+{\bf D}^{\top}$. Once these restricted transformation matrices have been constructed, full transformations satisfying the given constraint can be obtained by extending to the full space.
Remark 21
If $\hat{\bf S}$ is positive definite but nearly singular, then the coefficients of the Riccati inequality in (17) may incur large relative errors when formed. In this case, it may be appropriate to regularize the problem by perturbing $\hat{\bf S}$ to a nearby positive semidefinite (but now singular) problem as in (62), where the matrix ${\bf S}_{2}$ is now well conditioned with respect to inversion. In our procedure we do this by setting small positive eigenvalues of $\hat{\bf S}$ to zero.
@@ 4.2 Construction in the case ${\bf D}=-{\bf D}^{\top}$
Proof. The proof follows directly from (65) and (64) and the preceding results.
Note that without the constraint (66a), if $\bf A$} is stable, then by Lemma 9, the second condition (66b) can always be satisfied. Adding the constraint (66a), however, makes the question nontrivial.
@@
Lemma 23
[63] Consider ${\bf B},{\bf C}^{\top}\in{\mathbb{R}}^{n\times m}$, and assume that $\mathop{\rm rank}\nolimits{\bf B}=r$.
@@
Proof. Condition (66a) is equivalent to ${\bf C}={\bf B}^{\top}{\bf T}^{\top}{\bf T}$ and the following conditions will be necessary for the existence of an invertible $\bf T$} with this property:
@@
with ${\bf B}_{1},{\bf C}_{1}^{\top}\in{\mathbb{R}}^{n\times r}$ having full column rank. The conditions $\mathop{\rm rank}\nolimits{\bf C}{\bf B}=r$ and ${\bf C}{\bf B}\geq 0$ together are equivalent to ${\bf C}_{1}{\bf B}_{1}>0$. Thus, there must exist an invertible matrix $\bf Y$} (e.g., a Cholesky factor or a positive-definite square root) such that ${\bf C}_{1}{\bf B}_{1}={\bf Y}{\bf Y}^{\top}$.
@@
In order to explicitly construct a transformation matrix ${\bf T}_{0}$ as in (67), it will be useful to construct bi-orthogonal bases for the two subspaces $\mathsf{Ker}\,{\bf B}^{\top}$ and $\mathsf{Ker}\,{\bf C}$. Toward this end, let ${\bf N}_{{\bf C}}$ contain columns that form a basis of $\mathsf{Ker}\,{\bf C}$, so that $\mathsf{Ran}\,{\bf N}_{{\bf C}}=\mathsf{Ker}\,{\bf C}$. Such a matrix is easily constructed in a numerically stable way via the singular value decomposition or a rank-revealing $QR$ decomposition of $\bf C$}, see [25]. Since $\bf B$} and $\bf C$$^{\top}$ are assumed to satisfy $\mathsf{Ker}\,{\bf B}=\mathsf{Ker}\,{\bf C}^{\top}$, we have singular value decompositions
@@
Proof. The condition follows from Corollary 22 and the representation (67) by setting ${\bf U}={\bf I}$ and ${\bf T}_{0}$ as in (75).
Remark 25
It is clear that the matrices $\bf W$} and $\bf Y$} in Lemma 23 are not unique. In fact any matrix ${\bf W}=\left[\begin{array}[]{cc}{\bf W}_{11}&0\\ {\bf W}_{21}&{\bf W}_{22}\end{array}\right]$ with invertible ${\bf W}_{11}$ and ${\bf W}_{22}$ can replace $\bf W$}. As a consequence, ${\bf B}_{1}$ and ${\bf C}_{1}$ will be replaced by ${\bf B}_{1}{\bf W}_{11}$ and ${\bf W}_{11}^{\top}{\bf C}_{1}$. Then $\bf Y$} can be replaced by ${\bf W}_{11}^{\top}{\bf Y}{\bf U}_{\bf Y}$ for any real orthogonal matrix ${\bf U}_{\bf Y}$, and ${\bf Y}^{-1}{\bf C}_{1}$, ${\bf B}_{1}{\bf Y}^{-\top}$ become ${\bf U}_{\bf Y}^{\top}{\bf Y}^{-1}{\bf C}_{1}$ and ${\bf B}_{1}{\bf Y}^{-\top}{\bf U}_{\bf Y}$. The matrices $\widetilde{\bf N}_{\bf B}$ and $\widetilde{\bf N}_{\bf C}$ are not unique either. They can be replaced by $\widetilde{\bf N}_{\bf B}{\bf N}^{\top}$ and $\widetilde{\bf N}_{\bf C}{\bf N}^{-1}$ with an arbitrary invertible matrix $\bf N$}. In the end, ${\bf T}_{0}$ can be replaced by $\left[\begin{array}[]{cc}{\bf N}&0\\ 0&{\bf U}_{\bf Y}^{\top}\end{array}\right]{\bf T}_{0}$. The orthogonal matrix ${\bf U}_{\bf Y}$ can be absorbed in the matrix $\bf U$} in the formula for $\bf T$} in Lemma 23. The matrix $\bf Z$} in ${\bf T}_{\bf Z}$ needs to be replaced by ${\bf Z}{\bf N}^{-1}$. However, ${\bf Z}\widetilde{\bf N}_{\bf B}^{\top}$ is independent of $\bf N$}. Note that $\bf U$} affects $\bf T$} but not ${\bf Q}={\bf T}^{\top}{\bf T}$.
@@ 4.3 Construction in the case of general $\bf D$}
For the case that $\bf D$} is general we will present a recursive procedure which is analogous to the index reduction procedure for differential-algebraic equations in [38]. The first step is to perform the transformations (62), (63), and to obtain the following characterization when a transformation to port-Hamiltonian form (6) exists.
Lemma 26
Consider system (5) transformed as in (62) and (63). Then the system is equivalent to a port-Hamiltonian system of the form (6) if and only if the following two conditions hold
@@
    $\mathsf{Ker}\,{\bf C}_{1}^{\top}=\mathsf{Ker}\,{\bf B}_{1}$, $\mathop{\rm rank}\nolimits{\bf C}_{1}{\bf B}_{1}=\mathop{\rm rank}\nolimits{\bf B}_{1}{\color[rgb]{0,0,0}=r}$, and ${\bf C}_{1}{\bf B}_{1}\geq 0$, or equivalently, there exists an invertible (orthogonal) matrix $\bf W$} such that
@@
[…]
@@
    and the columns of full rank matrices $\widetilde{\bf N}_{{\bf B}_{1}}$ and $\widetilde{\bf N}_{{\bf C}_{1}}$ form the kernels of ${\bf B}_{1}^{\top}$, ${\bf C}_{1}$ respectively, satisfying $\widetilde{\bf N}_{{\bf B}_{1}}^{\top}\widetilde{\bf N}_{{\bf C}_{1}}={\bf I}_{n-r}$.
Proof. Transforming the system to the form (62) and (63), the inequality (64) can only hold if the second block column and row are zero. This gives the characterization in part 1., since in this case condition (66a) has the form $({\bf T}{\bf B}_{1})^{\top}={\bf C}_{1}{\bf T}^{-1}$. In this way the linear matrix inequality only has to be solved in the part where the feedthrough term has a positive definite symmetric part. The solvability for this part is then characterized by applying the result of Lemma 23 to ${\bf B}_{1}$ and ${\bf C}_{1}$ and one obtains $\bf T$} as the one in Lemma 23 b) with ${\bf T}_{0}$ of the form as (67). The result is then proved by applying this formula to (61).
In order to determine $\bf T$} we still need to find $\bf Z$}. Using the partition in the transformation matrix in (79) we obtain three blocks in $\widetilde{\bf Y}$ and we can repartition the middle factor as
$\left[\begin{array}[]{cc}-{\bf T}_{0}{\bf A}{\bf T}_{0}^{-1}&-{\bf T}_{0}{\bf B}_{2}\\ {\bf C}_{2}{\bf T}_{0}^{-1}&{{\bf D}_{2}}\end{array}\right]$ $=$ $\left[\begin{array}[]{c|cc}-\widetilde{\bf N}_{{\bf B}_{1}}^{\top}{\bf A}\widetilde{\bf N}_{{\bf C}_{1}}&-\widetilde{\bf N}_{{\bf B}_{1}}^{\top}{\bf A}\widehat{\bf B}_{1}{\bf Y}^{-\top}&-\widetilde{\bf N}_{{\bf B}_{1}}^{\top}{\bf B}_{2}\\ \hline\cr-{\bf Y}^{-1}\widehat{\bf C}_{1}{\bf A}\widetilde{\bf N}_{{\bf C}_{1}}&-{\bf Y}^{-1}\widehat{\bf C}_{1}{\bf A}\widehat{\bf B}_{1}{\bf Y}^{-\top}&-{\bf Y}^{-1}\widehat{\bf C}_{1}{\bf B}_{2}\\ {\bf C}_{2}\widetilde{\bf N}_{{\bf C}_{1}}&{\bf C}_{2}\widehat{\bf B}_{1}{\bf Y}^{-\top}&{{\bf D}_{2}}\end{array}\right]$ (86)
$=:$ $\left[\begin{array}[]{rr}-\widetilde{\bf A}_{1}&-\widetilde{\bf B}_{1}\\ \widetilde{\bf C}_{1}&\widetilde{\bf D}_{1}\end{array}\right].$ (89)
Thus, we have
$\widetilde{\bf Y}{\color[rgb]{0,0,0}({\bf Z})}=\left[\begin{array}[]{cc}{\bf Z}&0\\ 0&{{\bf I}_{r+s}}\end{array}\right]\left[\begin{array}[]{rr}-\widetilde{\bf A}_{1}&-\widetilde{\bf B}_{1}\\ \widetilde{\bf C}_{1}&\widetilde{\bf D}_{1}\end{array}\right]\left[\begin{array}[]{cc}{\bf Z}&0\\ 0&{{\bf I}_{r+s}}\end{array}\right]^{-1},$
@@
This implies that the procedure of checking the existence of a transformation from (5) to a pH system can be performed in a recursive way. One first performs the transformation (62) and checks whether a condition as in Part 1 of Lemma 26 holds. If this is not the case then the transformation to port-Hamiltonian form does not exist. If the condition holds, then one checks whether $\widetilde{\bf D}_{1}+\widetilde{\bf D}_{1}^{\top}$ in (89) is positive definite. If it is and if the associated matrix inequality does not have a positive definite solution, then a transformation to port-Hamiltonian form does not exist. Otherwise, a transformation matrix $\bf T$} can be constructed by computing $\bf Z$} satisfying $\widetilde{\bf Y}({\bf Z})+\widetilde{\bf Y}({\bf Z})^{\top}\geq 0$ and the matrix ${\bf T}_{\bf Z}=\left[\begin{array}[]{cc}{\bf Z}&0\\ 0&{\bf I}\end{array}\right]$ is formed accordingly. In the remaining case, i.e., if a condition as in Part 1 of Lemma 26 holds and $\widetilde{\bf D}_{1}+\widetilde{\bf D}_{1}^{\top}$ is singular, then the process is repeated with the reduced problem.
@@
[…]
@@
${\bf G}_{1}=\left[\begin{array}[]{c|cc|c}-\widetilde{\bf N}_{{\bf B}_{1}}^{\top}{\bf A}\widetilde{\bf N}_{{\bf C}_{1}}&-\widetilde{\bf N}_{{\bf B}_{1}}^{\top}{\bf A}\widehat{\bf B}_{1}{\bf Y}^{-\top}&-\widetilde{\bf N}_{{\bf B}_{1}}^{\top}{\bf B}_{2}&0\\ \hline\cr-{\bf Y}^{-1}\widehat{\bf C}_{1}{\bf A}\widetilde{\bf N}_{{\bf C}_{1}}&-{\bf Y}^{-1}\widehat{\bf C}_{1}{\bf A}\widehat{\bf B}_{1}{\bf Y}^{-\top}&-{\bf Y}^{-1}\widehat{\bf C}_{1}{\bf B}_{2}&-{\color[rgb]{0,0,0}\widetilde{\Gamma}}\\ {\bf C}_{2}\widetilde{\bf N}_{{\bf C}_{1}}&{\bf C}_{2}\widehat{\bf B}_{1}{\bf Y}^{-\top}&{\bf V}_{0,2}^{\top}{\bf D}{\bf V}_{0,2}&{\bf V}_{0,2}^{\top}{\bf D}{\bf V}_{0,1}\\ \hline\cr 0&{\color[rgb]{0,0,0}\widetilde{\Gamma}^{\top}}&{\bf V}_{0,1}^{\top}{\bf D}{\bf V}_{0,2}&{\bf V}_{0,1}^{\top}{\bf D}{\bf V}_{0,1}\end{array}\right],$
@@
${\bf C}_{1}{\bf T}_{0}^{-1}$ $=$ $\left[\begin{array}[]{cc}0&{\bf W}^{-\top}\left[\begin{array}[]{c}\widehat{\bf C}_{1}\\ 0\end{array}\right]\widehat{\bf B}_{1}{\bf Y}^{-\top}\end{array}\right]=\left[\begin{array}[]{cc}0&{\bf W}^{-\top}\left[\begin{array}[]{c}{\bf Y}\\ 0\end{array}\right]\end{array}\right]=\left[\begin{array}[]{cc}0&{\color[rgb]{0,0,0}\widetilde{\Gamma}^{\top}}\end{array}\right$],
${\bf T}_{0}{\bf B}_{1}$ $=$ $\left[\begin{array}[]{c}0\\ {\bf Y}^{-1}\widehat{\bf C}_{1}\left[\begin{array}[]{cc}\widehat{\bf B}_{1}&0\end{array}\right]{\bf W}^{-1}\end{array}\right]=\left[\begin{array}[]{c}0\\ \left[\begin{array}[]{cc}{\bf Y}^{\top}&0\end{array}\right]{\bf W}^{-1}\end{array}\right]=\left[\begin{array}[]{c}0\\ {\color[rgb]{0,0,0}\widetilde{\Gamma}}\end{array}\right$].